OpenStudy (anonymous):

Express the complex number in trigonometric form. 2i

OpenStudy (anonymous):



OpenStudy (anonymous):

from the picture you see that \(\theta=\frac{\pi}{2}\) and \(|2i|=2\) so it is \[2\left(\cos(\frac{\pi}{2})+i\sin(\frac{\pi}{2})\right)\]

OpenStudy (anonymous):

the check is easy, since \(\cos(\frac{\pi}{2})=0\) and \(\sin(\frac{\pi}{2})=1\) you get \(2i=2(0+i)=2i\)

OpenStudy (anonymous):

Okay so the degree of cos and isin would be 90